OpenStudy (anonymous):

23. During a clothing store’s Bargain Days, the regular price for T-shirts is discounted by $5. There is a state sales tax of 5%, and the $5 discount is applied before the sales tax is calculated.a. Write an expression that shows the regular price r of a T-shirt with the $5 discount applied.b. Write a rule for the function p(r) that expresses the final price p of a T-shirt with the discount applied and sales tax added.c. How much would you pay during Bargain Days for a shirt regularly priced at $15.50?

OpenStudy (anonymous):

a) d = r - 5 where t is the ful price b) p = ( r - 5 ) * 1.05 c) p = (15.5-5)*1.05 = 10.5 * 1.05 = $ 11.02

OpenStudy (mathstudent55):

Price of shirt: \(r\) Price including discount: \(r - 5\) \(p(r) = 1.05(r - 5)\) \(p(15.5) = 1.05(15.5 - 5) \)